The perimeter of a rhombus is 24 cm. The measure of its one diagonal is 8 cm. Find the area of rhombus.
Answer by addingup(3677)

Area = 35.78
---------------------------------
How?
A = 1/4d*sqrt(P^2-4d^2)where d=diagonal and P=perimeter
A = 1/4(8)*sqrt(24^2-4(8)^2)Use your calculator.